### v1

# Constrained Decoding for Robotics Foundation Models constrained-robot-fms.github.io

Abstract
Recent advances in the development of robotic foundation models have led to promising end-to-end and general-purpose capabilities in robotic systems. These models are pretrained on vast datasets of robot trajectories to process multimodal inputs and directly output a sequence of action that the system then executes in the real world. Although this approach is attractive from the perspective of improved generalization across diverse tasks, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness and safety constraints. We address these limitations by introducing a constrained decoding framework for robotics foundation models that enforces logical constraints on action trajectories in dynamical systems. Our method ensures that generated actions provably satisfy signal temporal logic (STL) specifications at runtime without retraining, while remaining agnostic of the underlying foundation model. We perform comprehensive evaluation of our approach across state-of-the-art navigation foundation models and we show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action-generation. Videos and code are available on our website: https://constrained-robot-fms.github.io
1 Introduction
Recent advances in Robotics Foundation Models (RFM) have enabled general purpose robot policies that map multi-modal inputs such as RGB images, natural language instructions and proprioceptive inputs to action sequences [1]. RFMs such as SPOC [2], PoliFormer [3], OpenVLA [4] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts. However, these models are primarily data-driven and lack any explicit notion of safety. Although these models may implicitly exhibit safety-related behaviors depending on the patterns in their training data, there is no formal guarantee that models will consistently behave safely in all situations. This serves as a limiting factor for deploying these foundation models in the physical world where rule compliance and regulatory safety rule adherence are crucial.
Formal specifications have long been used to specify rules of operation such as safety requirements or mission directives for robotic deployments [5, 6]. Specifically, temporal logics [7]  can capture rich desired temporal constraints over robot behavior, such as “remain within permitted region zones and avoid dangerous obstacles”. This provides a rigorous workflow to encode behavior rulebooks using temporal logics and enforce them for robotic foundation models. Although temporal logic has seen success in classical robotic planning for reliable behavior generation, its use for foundation models remains limited. Additionally, retraining or fine-tuning these large pre-trained models to directly embed temporal logic specification is challenging [8]. First, retraining models is costly endeavor in terms of computational resources and data requirements. Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of temporal constraints through training alone. Hence, there is a pressing need for methods that can enforce temporal specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time [9, 10, 11]. These approaches typically mask out tokens that violate a syntactic constraint defined over token sequences. For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns [9, 10]. Inspired by this line of work, we extend the paradigm of constrained decoding to enforce logical constraints over action trajectories in dynamical systems and propose specification aligned decoding (SpecDec) for RFMs that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [12] specifications. Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction. This conditioning is critical because it steers the model toward generating specification satisfying actions rather than relying on post hoc rejection. SpecDec reduces risk of infeasible outputs while preserving the original action distribution of the model. To enforce such specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations. Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to a lightweight dynamics model to predict future states. To efficiently evaluate STL specifications at inference time, we use a high-performance computational graph based library STLCG++ [13]. To the best of our knowledge, this is the first work to apply constrained decoding with STL specifications to RFMs.
Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for robotic foundation models under Signal Temporal Logic specifications (Section 3.1). Second, we propose an inference-time technique that reweights or masks candidate actions using STL satisfaction scores in (Section 3.2 and 3.3). Finally, we demonstrate the effectiveness of our method on state-of-the-art object navigation models without modifying model parameters (Section 4).
2 Preliminaries
2.1 Signal Temporal Logic
Signal Temporal Logic (STL) is an expressive framework for defining properties and reasoning over continuous time real valued signals [14]. Formally, $(\mathbf{s},t)\models\phi$ denotes that a signal $\mathbf{s}$ satisfies the STL formula $\phi$ at time $t$. An atomic predicate of an STL formula is represented by inequalities of the form $\mu(\mathbf{s}(t))>0$. The truth value of the predicate $\mu$ is equivalent to $\mu(\mathbf{s}(t))>0$. Note that with slight abuse of notation, $\mu$ represents both the predicate and a function of the trajectory $\mathbf{s}(t)$. Any STL formula consists of Boolean and temporal operations on these predicates, and the syntax of STL formulas is defined recursively as follows:
$\phi:=\mu~|~\neg\mu~|~\phi\land\psi~|~\phi\lor\psi~|~\mathbf{G}_{[a,b]}~\psi~|~\mathbf{F}_{[a,b]}~\psi~|~\phi~\mathbf{U}_{[a,b]}~\psi$
where $\psi$ and $\phi$ are STL formulae, $\mathbf{G}$ denotes the globally operator, $\mathbf{F}$ the eventually operator, and $\mathbf{U}$ is the until operator. For example, $\mathbf{s}\models\mathbf{G}_{[a,b]}\psi$ specifies that $\psi$ must be in all times in the given interval, $t\in[a,b]$ of the signal $\mathbf{s}$. Similarly, the operator until in $\mathbf{s}\models\phi\mathbf{U}_{[a,b]}\psi$ defines that $\phi$ must be true until $\psi$ becomes true within a time interval $[a,b]$.
Given a signal $s_{t}$ representing a signal starting at time t, the Boolean semantics of satisfaction of $s_{t}\models\phi$ are defined inductively as follows:
$s_{t}\models\mu$ $\iff\mu(s(t))>0$
$s_{t}\models\lnot\varphi$ $\iff\lnot(s_{t}\models\varphi)$
$s_{t}\models\varphi_{1}\land\varphi_{2}$ $\iff(s_{t}\models\varphi_{1})\land(s_{t}\models\varphi_{2})$
$s_{t}\models\text{F}_{[a,b]}(\varphi)$ $\iff\exists t^{\prime}\in[t+a,t+b]\text{ s.t. }s_{t^{\prime}}\models\varphi$
$s_{t}\models\text{G}_{[a,b]}(\varphi)$ $\iff\forall t^{\prime}\in[t+a,t+b]\text{ s.t. }s_{t^{\prime}}\models\varphi$
Apart from the Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric indicating robustness, i.e., the strength of satisfaction or violation. For the sake of brevity, the definition of robustness is provided in Appendix A.
2.2 Constrained Decoding in Transformers
A large variety of autoregressive transformer-based models generate final outputs by producing a probability distribution over the model vocabulary at each timestep. This distribution is generated by performing a softmax operation over the model’s last hidden layer. Then, through the process of decoding, tokens are selected to maximize the overall likelihood of an output sequence. In standard decoding, this maximization can be performed by either greedily selecting the most probable token at each step or by using a beam search to maintain multiple high-likelihood candidates. However, this often leads to degenerate output sequences that are repetitive [15]. A common approach is to use sampling strategies like top $k$ [16], and nucleus sampling [15] that introduce stochasticity to encourage more diverse outputs. Constrained decoding [17] modifies this probabilistic selection by pruning invalid tokens to ensure that the generated sequences satisfy predefined constraints. These constraints are often syntactic, such as regular expressions, JSON formatting, or programming language grammars [18]. There is also recent work on enforcing semantic constraints that ensure coherence of the output or alignment with specific knowledge bases [19]. Formally, constrained decoding can be seen as maximizing the probability of the output sequence subject to a constraint $\mathcal{C}$: $\arg\max_{y\in\mathcal{Y}_{\mathcal{C}}}P(y\mid x)$ where $\mathcal{Y}_{\mathcal{C}}$ is the set of sequences satisfying $\mathcal{C}$.
3 Specification-Guided Constrained Decoding
Figure 1: Overview of our specification aligned decoding framework. Given multimodal inputs (e.g., RGB images and natural language instructions), a pretrained Robot Foundation Model (e.g., SPOC) generates candidate actions. These actions are filtered or reweighted by the constrained decoding technique based on temporal logic constraints. In the figure, green markers denote target object locations, while red zones represent regions to avoid, as defined by temporal logic constraints.
In this section, we introduce a novel problem formulation for SpecDec in RFMs. First, we highlight the challenge in specification checking for RFMs in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it. Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
3.1 Problem Statement
As highlighted in the background section, existing techniques in constrained decoding for language models enforce syntactic constraints defined over tokens such as conforming to a context-free grammar or matching a regular expression. In these setups, constraint checking can be performed in the model’s token space.
In contrast, RFMs operate in a physical environment and constraints (captured via temporal logic) are defined over state variables. Since a large class of end-to-end RFMs solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward. In this case, constrain checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function. Hence, we leverage a dynamics function to compute specification satisfaction of different action sequences proposed by an RFM.
Consider a discrete dynamical system with states $x_{t}\in\mathbb{R}^{n}$ and actions $a_{t}\in\mathbb{A}$ at time step $t$. The system’s dynamics are defined by $x_{t+1}=f(x_{t},a_{t})$ where $f:\mathbb{R}^{n}\times\mathbb{A}\rightarrow\mathbb{R}^{n}$ maps the current state $(x_{t}\in\mathbb{R}^{n})$ and a discrete action ($a_{t}\in A)$ to the next state $x_{t+1}\in\mathbb{R}^{n}$. This system is controlled by a policy that selects an action $a_{t}$ at each time step based on observations and task context such as user-provided natural-language instructions or goal waypoints.
In this work, we focus on RFMs that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions. Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$. These inputs are first encoded into a latent embedding space through modality-specific encoders: $e_{\mathcal{I}_{t}}=\mathcal{E}_{\mathcal{I}}\!\bigl{(}\mathcal{I}_{t}\bigr{)}.$
Given the history of encoded inputs up to time $t$, a Transformer-based foundation model parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
$\{\hat{e}_{a_{t+k}}\}_{k=1}^{T-t}=\operatorname{Transformer}_{\theta}\!\bigl{(}\,\{e_{\mathcal{I}_{\tau}}\}_{\tau=0}^{t}\bigr{)}.$
Each predicted action embedding is decoded into an action $\hat{a}_{t+k}\in\mathbb{A}$, resulting in a predicted action sequence $\{\hat{a}_{t+1},\dots,\hat{a}_{T}\}$, where $T$ denotes the planning horizon.
Now, consider that the system is required to satisfy requirements encoded using an STL formula $\varphi$ defined over the state variables of the system. Formally, the goal is to ensure that the resulting trajectory satisfies the specification $\varphi$:
$\{(x_{0},\hat{a}_{0}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed foundation model actions through filtering or rejecting action sequences that violate the specification $\varphi$. Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low likelihood outputs. This undermines the inductive biases learned during pretraining and leads to degenerate, brittle behaviors. A similar problem was highlighted when ensuring compliance with logical constraints for large language models in [20]. Additionally, RFMs decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$. Posthoc manipulation can disrupt this causal chain and lead to a mismatch between the model’s internal hidden state and the executed sequence. Hence, we propose the following problem statement:
How can we enforce temporal logic constraints during action generation in robot foundation models such that the output sequence (1) satisfies an STL specification $\varphi$, and (2) remains faithful to the model’s autoregressive distribution $\pi(a_{1:T}\mid\mathcal{I}_{1:T})$?
Let $\pi(a_{1:T})$ be the unconstrained action-sequence distribution produced by the RFM’s decoder (e.g. the softmax over logits generated by the Transformer). We define the ideal constrained distribution over action sequences as:
$Q_{\pi,\varphi}(a_{1:T})=\frac{\pi(a_{1:T})\cdot\mathbf{1}[(x_{1:T},a_{1:T})\models\varphi]}{\sum\limits_{a_{1:T}^{\prime}}\pi(a_{1:T}^{\prime})\cdot\mathbf{1}[(x_{1:T}^{\prime},a_{1:T}^{\prime})\models\varphi]}$ (1)
where $x_{1:T}$ denotes the state trajectory induced by the system dynamics under actions $a_{1:T}$ and $\mathbf{1}[\cdot]$ is the indicator function that returns 1 iff the trajectory-action pair satisfies the specification. Equation 1 is the exact Bayesian conditioning of $\pi$ on the event that the generated rollout satisfies $\varphi$. Hence, sampling from $Q_{\pi,\varphi}$ would give sequences that (i) inherit the original model’s preferences encoded in $\pi$ and (ii) guarantee specification satisfaction.
In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques. Specifically, we propose SpecDec: A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
3.2 Hard Constrained Decoding
As highlighted in the background section, in the final layer, predictions are detokenized and a projection layer converts the embeddings into logits over the vocabulary space. These logits are further converted into a probability distribution using a softmax operation. In prior work, for structured output generation in LLMs, some invalid tokens are masked based on syntactical constraints or other criteria [18, 20]. This is done by setting their logit value as -$\infty$ before the softmax operation is applied. For HCD, we use a similar approach as constrained decoding literature [18] and mask out predicted action tokens that violate our given STL specification $\varphi$ during sequential generation. Formally, to enforce the STL specification $\varphi$ during sequential generation, we adjust the logits at each timestep $t+k$ as follows:
Let $\textbf{z}_{t+k}$ denote the logits at timestep $t+k$. For each action choice $i$ at timestep $t+k$, we define:
$z_{t+k}^{(i)}=\begin{cases}-\infty,&\text{if }\hat{x}_{t+k}^{(i)}=f(x_{t+k-1},\hat{a}_{t+k}^{(i)})\text{ violates }\varphi\\ z_{t+k}^{(i)},&\text{otherwise}\end{cases}$
Here $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action. This next state is elicited using a simple dynamics model ($f$) as highlighted in the previous section. Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function. We introduce a theorem to state that our proposed approach generates a trajectory that satisfies the given specification $\varphi$:
Theorem 1 (Specification Satisfaction Modulo Dynamics).
Let $\pi_{\text{HCD}}$ denote the policy induced by hard-constrained decoding with respect to an STL specification $\varphi$, and let $f$ be a deterministic dynamics model used during decoding. Then the generated trajectory satisfies the specification:
$\{(x_{t+1},\hat{a}_{t+1}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
where each action $\hat{a}_{t+k}\sim\pi_{\text{HCD}}(\cdot\mid x_{t+k-1})$, and $x_{t+k}=f(x_{t+k-1},\hat{a}_{t+k})\quad\text{for all }k.$
A proof sketch is provided in the Appendix B
3.3 Robustness Constrained Decoding
HCD ensures compliance but can lead to compromising task success, which can be undesirable. A similar tradeoff was observed by [21] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity. Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness). Unlike HCD, which applies hard masking to completely remove unsafe actions, RCD softly guides the model toward safer actions by incorporating robustness scores that reflect the degree of satisfaction of $\varphi$. This is similar to the approach proposed in [21] where the next-token distributions were re-weighted based on the utility scores provided by another language model. Our utility scores are quantified by the robustness function( $\rho(x_{t},\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification. Positive robustness values denote specification satisfaction, while negative values capture the degree of violation.
First, we compute a robustness score for each candidate action: $r_{t+k}^{(i)}=\rho(\hat{x}_{t+k}^{(i)},\varphi)$ where $\rho(\cdot,\varphi)$ is the STL robustness metric, and $\hat{x}_{t+k}^{(i)}$ is the predicted next state under action $\hat{a}_{t+k}^{(i)}$. This robustness score $r_{t+k}^{(i)}$ quantifies how well each candidate action satisfies the specification $\varphi$. These scores are then converted into weights using exponential scaling: $w_{t+k,i}=\exp(\alpha\cdot r_{t+k,i})$ where $\alpha$ is a temperature parameter that adjusts the sharpness of the bias. We use these weights to shift the original logits: $\tilde{z}_{t+k,i}=z_{t+k,i}+\beta\cdot w_{t+k,i}$ where $\beta$ is a hyperparameter that modulates the trade-off between specification adherence and the original task objective. Finally, we obtain the action distribution by applying softmax over the adjusted logits: $p_{t+k}=\text{softmax}(\tilde{\mathbf{z}}_{t+k})$
This approach allows for graded preferences that improve flexibility and robustness to dynamics approximation errors. This is because every action keeps a non-zero probability, the policy can still recover when the dynamics model is imperfect. Concretely, if the predicted successor $\hat{x}_{t+1}$ is off by $\epsilon$, an action that looked marginally unsafe can be safe in the true system, and vice versa. Retaining a weighted down probability for this action gives the sampler a fall-back option where as HCD would completely rule this action out due to 0 probability. Since we are shifting the probability mass for unsatisfying actions, it is possible that they are still chosen and lead to a violation. However, this is a tradeoff we allow to achieve a given task objective. We note that this still ensures higher STL sastifaction than unconstrained actions.
4 Evaluation
4.1 Implementational Details
We evaluate our constrained decoding approach on AI2-THOR simulated [22] indoor environments using the state-of-the-art (SOTA) navigation model called Shortest Path Oracle Clone (SPOC) [2]. This model achieves high navigation success rates for diverse tasks such as object navigation, room navigation, and so forth. We encode the STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [13]. This ensures minimal inference overhead at runtime, which is crucial for foundation model deployment. We enforce geofencing and obstacle avoidance by encoding them into invariant STL safety specifications. Specifically, we generate random regions in the configuration space that the robot must either avoid (obstacle zones) or remain within (safe zones), and apply these constraints in real time during execution.
4.2 Benchmarks and Research Questions
We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism. The filtering mechanism picks a default action (turning left or right in place) upon predicted violation of the safety specification, similar to the Simplex architecture [23]. Simplex architecture is a classic scheme in which a high-performance advanced controller is continuously monitored by a provably safe but less capable backup controller. Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons. We are interested in three main metrics: STL Satisfaction Rate (STL St): Proportion of trajectories that satisfy the specified STL formula, Task Success rate (SR): Standard task success, Average Robustness Score (AR): A quantitative measure of how strongly STL constraints are satisfied. The three main research questions we investigated in this paper:
  1. 1.
    RQ1: Do RCD and HCD outperform the baselines in STL satisfaction?
  2. 2.
    RQ2: Do RCD and HCD lead to task success rates comparable to those of the unconstrained technique?
  3. 3.
    RQ3: Does RCD perform better than HCD for task success rates?
Figure 2: Distribution of STL robustness scores across three action selection techniques for $\phi_{avoid}$. The red dashed line marks the violation threshold at robustness = 0. HCD achieves the lowest violation rate, while RCD strikes a balance between task success and STL satisfaction.
Specification Decoding STL St (%) SR (%) AR
$\phi_{avoid}$ Unconstrained (SPOC) 72.00 82.5 1.60
Filtering (SPOC-F) 97.00 72.00 1.57
Hard-Constrained (SPOC-HCD) 97.00 72.50 1.73
Robust-Constrained (SPOC-RCD) 93.00 76.00 1.71
$\phi_{geofence}$ Unconstrained (SPOC) 78.00 81.5 1.68
Filtering (SPOC-F) 100.00 72.00 1.89
Hard-Constrained (SPOC-HCD) 100.00 76.5 1.97
Robust-Constrained (SPOC-RCD) 95.35 80.23 1.86
Table 1: Comparison of decoding strategies across different backbone models on STL-constrained tasks. Our constrained decoding approaches show significant improvements in STL satisfaction and robustness while preserving task success.
4.3 Results
Our results are highlighted in Table 1 and Figure 2. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
RQ1 – STL satisfaction.
[…]
RQ2 – Task completion.
[…]
RQ3 – RCD vs. HCD.
While both HCD and RCD improve safety over the unconstrained baseline, they differ in how they balance constraint satisfaction with task success. HCD enforces strict STL satisfaction that results in frequent conservatism and lower successful task completion rates. In contrast, RCD’s soft penalization leads to higher task success—up to 4–5 % more than HCD—while still maintaining over 93 % STL satisfaction. These results show that RCD achieves a better trade-off between safety and goal-directed behavior, especially in settings where occasional low-risk actions can lead to higher long-term rewards. We also perform an ablation for varying values of $\beta$ to investigate this tradeoff and provide these results in Appendix C.
Our proposed techniques effectively enforces STL specifications during policy execution. HCD ensures full compliance but occasionally sacrifices task success due to strict truncation. RCD strikes a balance, offering high satisfaction rates and robust performance. This highlights the feasibility of combining learning-based models with formal safety constraints.
5 Related Work
Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [24], safe reinforcement learning [25], and temporal logic-based shielding approaches [26]. Recently, with the advent of Vision Language Action models and their impressive generalizable capabilities for manipulation, navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models. Although classical methods offer formal guarantees, they require pretraining/fine-tuning stage interventions, which can be restrictive. For example, SafeVLA [27] fine-tunes pre-trained foundation models with task-specific safety costs, achieving strong performance in Safety-CHORES tasks. However, the safety specification is expected to be embedded in the training data and loss, meaning the model cannot generalize to new safety constraints at test time. In contrast, ASIMOV [28] explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees. Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model and incurring negligible computational overhead. The closest to our work is SELP [29] that proposes LTL-constrained decoding for language model-based plan generation. However, since LTL does not possess quantiative semantics, they are restricted to hard masking-based approach.
6 Limitations and Future Work
In this work, we introduce a constrained decoding framework for enforcing STL specifications for robotic foundation models. Our approach enables runtime adaptation to novel safety specifications without retraining. Through experiments across multiple simulated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates. Our approach makes two critical assumptions that can be a limiting factor. First, we assume access to specifications that are defined over the state space and that these specifications are generated by roboticists. Although this is a common situation for safety critical deployment contexts like aerial robotics [30, 31], these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation. We hope to remedy this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic (ETL) [32] and using Large Language Models for generating high level specifications automatically [33]. Second, our approach also assumes access to a predictive model to evaluate the impact of actions on future trajectories. This can limit applicability to settings where accurate dynamics models are unavailable. However, it is possible to mitigate this via learned dynamics models [34] or world models proposed in [35, 36]. In addition, while our experiments are conducted in simulation, the underlying foundation model (SPOC) has been successfully deployed on real robots with minimal sim-to-real degradation [2]. Since our constrained decoding operates primarily at inference time and leaves the base model unchanged, the observed improvements in STL satisfaction and task success are expected to transfer over to the real world. We plan to rigorously test this by deploying SpecDec on the same physical robot platform as [2].
Appendix
Appendix A Quantiative Semantics of STL
Given a signal $s_{t}$ representing a signal starting at time t, the quantitative semantics of satisfaction of $s_{t}\models\phi$ are defined inductively as follows:
$\rho(s_{t},\mu_{c})$ $=\mu(x_{t})-c$
$\rho(s_{t},\lnot\varphi)$ $=-\rho(s_{t},\varphi)$
$\rho(s_{t},\varphi_{1}\land\varphi_{2})$ $=\min(\rho(s_{t},\varphi_{1}),\rho(s_{t},\varphi_{2}))$
$\rho(s_{t},\text{F}_{[a,b]}(\varphi))$ $=\underset{t^{\prime}\in[t+a,t+b]}{\max}\rho(s_{t}^{\prime},\varphi)$
$\rho(s_{t},\text{G}_{[a,b]}(\varphi))$ $=\underset{t^{\prime}\in[t+a,t+b]}{\min}\rho(s_{t}^{\prime},\varphi)$
Appendix B HCD proof
We rewrite the theorem here for easier comprehension.
Let $\pi_{\text{HCD}}$ denote the policy induced by hard-constrained decoding with respect to an STL specification $\varphi$, and let $f$ be a deterministic dynamics model used during decoding. Then the generated trajectory satisfies the specification:
$\{(x_{t+1},\hat{a}_{t+1}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
where each action $\hat{a}_{t+k}\sim\pi_{\text{HCD}}(\cdot\mid x_{t+k-1})$, and $x_{t+k}=f(x_{t+k-1},\hat{a}_{t+k})\quad\text{for all }k.$
Proof Sketch.
[…]
We proceed by induction over time. At step $t+1$, the masking ensures that the selected action $a_{t+1}$ is such that $(x_{t+1},a_{t+1})\models\varphi$, where $x_{t+1}=f(x_{t},a_{t+1})$. Assuming that the sequence $\{(x_{t+1},a_{t+1}),\dots,(x_{t+k-1},a_{t+k-1})\}$ satisfies $\varphi$, we show that $(x_{t+k},a_{t+k})$ will also satisfy $\varphi$. Since actions are masked based on predicted rollouts using $f$, and $f$ is exact, the only permitted actions at step $t+k$ are those that extend the sequence without violating $\varphi$. By induction, the entire sequence satisfies $\varphi$. ∎
Appendix C Robustness weighting ablation
Figure 3: Success rate and STL satisfaction rate across different values of $\beta$ for RCD, which controls the importance of STL spcification during decoding. As $\beta$ increases, STL satisfaction improves while task success shows a slight tradeoff beyond moderate values.
To evaluate the impact of relative weighting between robustness and base logits, we plotted the success rate and STL satisfaction rate across varying values for $\beta$. As shown in Figure 3, increasing the robustness threshold initially improves both task success and specification satisfaction, with success rates peaking around a moderate threshold (5–10). However, at higher thresholds (20 and 25), success rates begin to decline slightly, while STL satisfaction remains high.
This trend highlights a trade-off introduced by weighting robustness. Higher robustness weighting leads to better STL satisfaction at the cost of hindering overall task completion as logits for feasible actions are shifted heavily by the robustness values. These results demonstrate the importance of carefully balancing robustness weighting to maintain both task success and STL satisfaction.

### v3

@@
-Constrained Decoding for Robotics Foundation Models constrained-robot-fms.github.io
+Constrained Decoding for Robotics Foundation Models
 Abstract
 Recent advances in the development of robotic foundation models have led to promising end-to-end and general-purpose capabilities in robotic systems.
-These models are pretrained on vast datasets of robot trajectories to process multimodal inputs and directly output a sequence of action that the system then executes in the real world.
-Although this approach is attractive from the perspective of improved generalization across diverse tasks, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness and safety constraints.
-We address these limitations by introducing a constrained decoding framework for robotics foundation models that enforces logical constraints on action trajectories in dynamical systems.
-Our method ensures that generated actions provably satisfy signal temporal logic (STL) specifications at runtime without retraining, while remaining agnostic of the underlying foundation model.
-We perform comprehensive evaluation of our approach across state-of-the-art navigation foundation models and we show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action-generation.
-Videos and code are available on our website: https://constrained-robot-fms.github.io
+Trained on vast datasets of simulated and real-world trajectories, these models map multimodal observations directly to action sequences for physical execution.
+Despite promising real-world capabilities, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness.
+We address this gap by introducing SafeDec, a constrained decoding framework for autoregressive, robot foundation models that enforces invariant safety specifications on candidate action trajectories.
+Task-specific safety rules are expressed as Signal Temporal Logic (STL) formulas and are enforced at inference time with minimal overhead.
+Our method ensures that generated actions provably satisfy STL specifications under assumed dynamics at runtime without retraining , while remaining agnostic of the underlying policy.
+We evaluate SafeDec on tasks from the CHORES benchmark for state-of-the-art generalist policies (e.g., SPOC, Flare, PoliFormer) across hundreds of procedurally generated environments and show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action generation.
+Videos are available at constrained-robot-fms.github.io.
 1 Introduction
-Recent advances in Robotics Foundation Models (RFM) have enabled general purpose robot policies that map multi-modal inputs such as RGB images, natural language instructions and proprioceptive inputs to action sequences [1].
-RFMs such as SPOC [2], PoliFormer [3], OpenVLA [4] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts.
+Recent advances in developing Robotics Foundation Models (RFM) have enabled general-purpose policies that map multi-modal inputs such as RGB images, natural language instructions, and proprioceptive inputs to action sequences Hu et al.
+[1].
+Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3], Flare [4] and OpenVLA [5] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts.
 However, these models are primarily data-driven and lack any explicit notion of safety.
@@
 This serves as a limiting factor for deploying these foundation models in the physical world where rule compliance and regulatory safety rule adherence are crucial.
-Formal specifications have long been used to specify rules of operation such as safety requirements or mission directives for robotic deployments [5, 6].
-Specifically, temporal logics [7]  can capture rich desired temporal constraints over robot behavior, such as “remain within permitted region zones and avoid dangerous obstacles”.
-This provides a rigorous workflow to encode behavior rulebooks using temporal logics and enforce them for robotic foundation models.
-Although temporal logic has seen success in classical robotic planning for reliable behavior generation, its use for foundation models remains limited.
+Formal specifications have long been used to specify safety requirements for robotic deployments [6, 7].
+Temporal logics (TL) [8]  can capture safety constraints on robot behavior, such as “remain within the permitted region zones and avoid dangerous obstacles”.
+Although TL has seen success in classical robotic planning for safety constraint satisfaction, its use for enforcing safety for large transformer-based robot policies remains limited.
 Additionally, retraining or fine-tuning these large pre-trained models to directly embed temporal logic specification is challenging [9].
-First, retraining models is costly endeavor in terms of computational resources and data requirements.
-Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of temporal constraints through training alone.
-Hence, there is a pressing need for methods that can enforce temporal specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
+First, retraining models is a costly endeavor in terms of computational resources and data requirements.
+Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of safety constraints through training alone.
+Hence, there is a pressing need for methods that can enforce safety specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
 In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time [10, 11, 12].
@@
 For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns [10, 11].
-Inspired by this line of work, we extend the paradigm of constrained decoding to enforce logical constraints over action trajectories in dynamical systems and propose specification aligned decoding (SpecDec) for RFMs that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [12] specifications.
+Inspired by this line of work, we extend the paradigm of constrained decoding to enforce safety constraints over action trajectories in dynamical systems and propose safety specification aligned decoding (SafeDec) for transformer based policies that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [13] specifications under assumed dynamics.
 Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction.
-This conditioning is critical because it steers the model toward generating specification satisfying actions rather than relying on post hoc rejection.
-SpecDec reduces risk of infeasible outputs while preserving the original action distribution of the model.
+This conditioning is critical because it steers the model toward generating safety specification satisfying actions rather than relying on post hoc rejection.
+SafeDec reduces risk of infeasible outputs while preserving the original action distribution of the model.
 To enforce such specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations.
-Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to a lightweight dynamics model to predict future states.
+Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to an approximate dynamics model to predict future states.
 To efficiently evaluate STL specifications at inference time, we use a high-performance computational graph based library STLCG++ [14].
-To the best of our knowledge, this is the first work to apply constrained decoding with STL specifications to RFMs.
-Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for robotic foundation models under Signal Temporal Logic specifications (Section 3.1).
+To the best of our knowledge, this is the first work to enforce formal safety on transformer-based robotic policies at inference time using constrained decoding.
+Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for transformer-based policies under safety STL specifications (Section 3.1).
 Second, we propose an inference-time technique that reweights or masks candidate actions using STL satisfaction scores in (Section 3.2 and 3.3).
@@
 Apart from the Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric indicating robustness, i.e., the strength of satisfaction or violation.
-For the sake of brevity, the definition of robustness is provided in Appendix A.
+For the sake of brevity, the definition of robustness is provided in Appendix A.1.
 2.2 Constrained Decoding in Transformers
@@
 Given multimodal inputs (e.g., RGB images and natural language instructions), a pretrained robot foundation model (e.g., SPOC) generates candidate actions.
-These actions are filtered or reweighted by the constrained decoding technique based on temporal logic constraints.
-In the figure, green markers denote target object locations, while red zones represent regions to avoid, as defined by temporal logic constraints.
-In this section, we introduce a novel problem formulation for SpecDec in RFMs.
-First, we highlight the challenge in specification checking for RFMs in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it.
+These actions are filtered or reweighted by the constrained decoding technique based on STL safety constraints.
+In the figure, green markers denote target object locations, while red zones represent regions to avoid.
+In this section, we introduce a novel problem formulation for SafeDec in RFMs.
+First, we highlight the challenge in specification checking for these models in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it.
 Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
 3.1 Problem Statement
-As highlighted in the background section, existing techniques in constrained decoding for language models enforce syntactic constraints defined over tokens such as conforming to a context-free grammar or matching a regular expression.
+As highlighted in the background section, existing techniques in constrained decoding for language models enforce syntactic constraints defined over tokens, such as conforming to a context-free grammar or matching a regular expression.
 In these setups, constraint checking can be performed in the model’s token space.
 In contrast, RFMs operate in a physical environment and constraints (captured via temporal logic) are defined over state variables.
-Since a large class of end-to-end RFMs solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward.
-In this case, constrain checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function.
-Hence, we leverage a dynamics function to compute specification satisfaction of different action sequences proposed by an RFM.
+Since a large class of end-to-end policies solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward.
+In this case, constraint checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function.
+Hence, we leverage an approximate first order dynamics function to compute specification satisfaction of different action sequences proposed by these policies.
 Consider a discrete dynamical system with states $x_{t}\in\mathbb{R}^{n}$ and actions $a_{t}\in\mathbb{A}$ at time step $t$.
@@
 Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$.
-These inputs are first encoded into a latent embedding space through modality-specific encoders: $e_{\mathcal{I}_{t}}=\mathcal{E}_{\mathcal{I}}\!\bigl{(}\mathcal{I}_{t}\bigr{)}.$
-Given the history of encoded inputs up to time $t$, a Transformer-based foundation model parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
+Given the history of encoded inputs up to time $t$, a transformer-based policy parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
 $\{\hat{e}_{a_{t+k}}\}_{k=1}^{T-t}=\operatorname{Transformer}_{\theta}\!\bigl(\,\{e_{\mathcal{I}_{\tau}}\}_{\tau=0}^{t}\bigr).$
@@
 $\{(x_{0},\hat{a}_{0}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
-Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed foundation model actions through filtering or rejecting action sequences that violate the specification $\varphi$.
-Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low likelihood outputs.
+Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed actions through filtering or rejecting action sequences that violate the specification $\varphi$.
+Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low-likelihood output.
 This undermines the inductive biases learned during pretraining and leads to degenerate, brittle behaviors.
-A similar problem was highlighted when ensuring compliance with logical constraints for large language models in [20].
-Additionally, RFMs decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$.
+A similar problem was highlighted when ensuring compliance with logical constraints for large language models (LLM) in Park et al.
+[21].
+Additionally, these models decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$.
 Posthoc manipulation can disrupt this causal chain and lead to a mismatch between the model’s internal hidden state and the executed sequence.
@@
 How can we enforce temporal logic constraints during action generation in robot foundation models such that the output sequence (1) satisfies an STL specification $\varphi$, and (2) remains faithful to the model’s autoregressive distribution $\pi(a_{1:T}\mid\mathcal{I}_{1:T})$?
-Let $\pi(a_{1:T})$ be the unconstrained action-sequence distribution produced by the RFM’s decoder (e.g.
+Let $\pi(a_{0:T})$ be the unconstrained action-sequence distribution produced by the transformer-based policy’s decoder (e.g.
 the softmax over logits generated by the Transformer).
 We define the ideal constrained distribution over action sequences as:
-$Q_{\pi,\varphi}(a_{1:T})=\frac{\pi(a_{1:T})\cdot\mathbf{1}[(x_{1:T},a_{1:T})\models\varphi]}{\sum\limits_{a_{1:T}^{\prime}}\pi(a_{1:T}^{\prime})\cdot\mathbf{1}[(x_{1:T}^{\prime},a_{1:T}^{\prime})\models\varphi]}$ (1)
-where $x_{1:T}$ denotes the state trajectory induced by the system dynamics under actions $a_{1:T}$ and $\mathbf{1}[\cdot]$ is the indicator function that returns 1 iff the trajectory-action pair satisfies the specification.
+$Q_{\pi,\varphi}(a_{0:T})=\frac{\pi(a_{0:T})\cdot\mathbf{1}[(x_{1:T},a_{0:T})\models\varphi]}{\sum\limits_{a_{0:T}^{\prime}}\pi(a_{0:T}^{\prime})\cdot\mathbf{1}[(x_{1:T}^{\prime},a_{0:T}^{\prime})\models\varphi]}$ (1)
+where $x_{1:T}$ denotes the state trajectory induced by the system dynamics under actions $a_{0:T}$ and $\mathbf{1}[\cdot]$ is the indicator function that returns 1 iff the trajectory-action pair satisfies the specification.
 Equation 1 is the exact Bayesian conditioning of $\pi$ on the event that the generated rollout satisfies $\varphi$.
@@
 In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques.
-Specifically, we propose SpecDec: A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
+Specifically, we propose SafeDec : A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
 3.2 Hard Constrained Decoding
@@
 Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function.
-We introduce a theorem to state that our proposed approach generates a trajectory that satisfies the given specification $\varphi$:
-Theorem 1 (Specification Satisfaction Modulo Dynamics).
-Let $\pi_{\text{HCD}}$ denote the policy induced by hard-constrained decoding with respect to an STL specification $\varphi$, and let $f$ be a deterministic dynamics model used during decoding.
-Then the generated trajectory satisfies the specification:
-$\{(x_{t+1},\hat{a}_{t+1}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
-where each action $\hat{a}_{t+k}\sim\pi_{\text{HCD}}(\cdot\mid x_{t+k-1})$, and $x_{t+k}=f(x_{t+k-1},\hat{a}_{t+k})\quad\text{for all }k.$
-A proof sketch is provided in the Appendix B
 3.3 Robustness Constrained Decoding
 HCD ensures compliance but can lead to compromising task success, which can be undesirable.
-A similar tradeoff was observed by [21] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity.
+A similar tradeoff was observed by Liu et al.
+[22] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity.
 Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness).
 Unlike HCD, which applies hard masking to completely remove unsafe actions, RCD softly guides the model toward safer actions by incorporating robustness scores that reflect the degree of satisfaction of $\varphi$.
-This is similar to the approach proposed in [21] where the next-token distributions were re-weighted based on the utility scores provided by another language model.
+This is similar to the approach proposed in Liu et al.
+[22] where the next-token distributions were re-weighted based on the utility scores provided by another language model.
 Our utility scores are quantified by the robustness function( $\rho(x_{t},\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification.
 Positive robustness values denote specification satisfaction, while negative values capture the degree of violation.
+A formal definition of robustness in line with the STL quantitative semantics is provided in Appendix A.1.
 First, we compute a robustness score for each candidate action: $r_{t+k}^{(i)}=\rho(\hat{x}_{t+k}^{(i)},\varphi)$ where $\rho(\cdot,\varphi)$ is the STL robustness metric, and $\hat{x}_{t+k}^{(i)}$ is the predicted next state under action $\hat{a}_{t+k}^{(i)}$.
@@
 This approach allows for graded preferences that improve flexibility and robustness to dynamics approximation errors.
-This is because every action keeps a non-zero probability, the policy can still recover when the dynamics model is imperfect.
-Concretely, if the predicted successor $\hat{x}_{t+1}$ is off by $\epsilon$, an action that looked marginally unsafe can be safe in the true system, and vice versa.
+By ensuring that all actions preserve a non-zero probability, the policy remains capable of recovering from errors arising from an imperfect dynamics model.
+Concretely, if the predicted successor $\hat{x}_{t+1}$ is off by $\epsilon$, an action that appeared marginally unsafe can be safe in the true system, and vice versa.
 Retaining a weighted down probability for this action gives the sampler a fall-back option whereas HCD would completely rule this action out due to 0 probability.
 Since we are shifting the probability mass for unsatisfying actions, it is possible that they are still chosen and lead to a violation.
-However, this is a tradeoff we allow to achieve a given task objective.
-We note that this still ensures higher STL sastifaction than unconstrained actions.
+However, this is a tradeoff that we allow to achieve a given task objective.
+We note that this still ensures higher STL satisfaction than unconstrained actions.
 4 Evaluation
 4.1 Implementational Details
-We evaluate our constrained decoding approach on AI2-THOR simulated [22] indoor environments using the state-of-the-art (SOTA) navigation model called Shortest Path Oracle Clone (SPOC) [2].
-This model achieves high navigation success rates for diverse tasks such as object navigation, room navigation, and so forth.
-We encode the STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [13].
-This ensures minimal inference overhead at runtime, which is crucial for foundation model deployment.
-We enforce geofencing and obstacle avoidance by encoding them into invariant STL safety specifications.
+We comprehensively evaluate our constrained decoding framework on procedurally generated AI2-THOR [23] indoor scenes with diverse objects and layouts using three state-of-the-art (SOTA) RFMs: Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3] and Flare [4].
+All three are large transformer-based embodied agents trained on extensive language-conditioned robot trajectory datasets.
+These models achieve strong zero-shot generalization for a vast variety of navigation tasks that span open vocabulary object-goal navigation (“find a mug”), room-to-room traversal (“visit all rooms”), waypoint-based navigation (“move three meters forward and stop near the red rug”), and attribute-conditioned variants (“locate the chair closest to the refrigerator in the kitchen”).
+These models also demonstrate reliable zero-shot transfer to real-world environments, achieving robust task satisfaction.
+In addition, these models capture three different training paradigms for RFMs.
+SPOC is trained purely with imitation learning from shortest-path rollouts.
+Poliformer employs a hybrid approach that combines reinforcement learning and imitation learning, enabling it to learn long-horizon structure while retaining expert priors.
+Flare adopts a large-scale pretraining plus fine-tuning on embodied navigation data in line with recent foundation model training paradigms.
+This diversity in training paradigms allows us to evaluate the applicability of SafeDec across different learning regimes.
+In this work, we address safety specifications for robotics and, therefore, select those most relevant to real-world deployment.
+In particular, we focus on an important class of safety specifications called invariants, which are specifications that must be enforced at every reachable state of the system (e.g., “always avoid an unsafe region”).
+We enforce geofencing and obstacle avoidance by encoding them as invariant specifications in STL.
 Specifically, we generate random regions in the configuration space that the robot must either avoid (obstacle zones) or remain within (safe zones), and apply these constraints in real time during execution.
-4.2 Benchmarks and Research Questions
+The specifications used are: $\varphi_{\text{geofence}}=\mathbf{G}\!\left(\bigvee_{i=1}^{N}\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$, $\varphi_{\text{avoid}}\;=\;\mathbf{G}\!\left(\bigwedge_{i=1}^{N}\neg\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$.
+The size of the regions for $\varphi_{avoid}$ is 1 $m^{2}$.
+For $\varphi_{geofence}$, we randomly pick a subset of rooms in each house and use each chosen room’s full bounds.
+These safety specifications are common for robot learning applications [24, 25, 26, 27].
+We encode our test STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [14].
+This ensures minimal inference overhead at runtime ($10^{-5}$ s per timestep) , which is crucial for policy deployment.
+For our dynamics model, we assume a unicycle model, an approximate first-order dynamics abstraction widely used in the robotics literature for analysis and control [28].
+This representation captures the essential kinematics of motion in the plane and is widely used because it is applicable for diverse robotic platforms.
+4.2 Experimental setup
 We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism.
@@
 Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons.
-We are interested in three main metrics: STL Satisfaction Rate (STL St): Proportion of trajectories that satisfy the specified STL formula, Task Success rate (SR): Standard task success, Average Robustness Score (AR): A quantitative measure of how strongly STL constraints are satisfied.
+We evaluate performance using two main metrics: STL Satisfaction Rate (STL St), defined as the proportion of trajectories that satisfy the specified STL formula, and Task Success Rate (SR), which measures standard task success.
 The three main research questions we investigated in this paper:
   1. 1.
-    RQ1: Do RCD and HCD outperform the baselines in STL satisfaction?
+    RQ1: Do HCD and RCD provide higher STL satisfaction than the unconstrained baselines?
   2. 2.
-    RQ2: Do RCD and HCD lead to task success rates comparable to those of the unconstrained technique?
+    RQ2: Do HCD and RCD preserve task success rates comparable to the unconstrained baselines?
   3. 3.
-    RQ3: Does RCD perform better than HCD for task success rates?
-Figure 2: Distribution of STL robustness scores across three action selection techniques for $\phi_{avoid}$.
-The red dashed line marks the violation threshold at robustness = 0.
-HCD achieves the lowest violation rate, while RCD strikes a balance between task success and STL satisfaction.
-Specification Decoding STL St (%) SR (%) AR
-$\phi_{avoid}$ Unconstrained (SPOC) 72.00 82.5 1.60
-Filtering (SPOC-F) 97.00 72.00 1.57
-Hard-Constrained (SPOC-HCD) 97.00 72.50 1.73
-Robust-Constrained (SPOC-RCD) 93.00 76.00 1.71
-$\phi_{geofence}$ Unconstrained (SPOC) 78.00 81.5 1.68
-Filtering (SPOC-F) 100.00 72.00 1.89
-Hard-Constrained (SPOC-HCD) 100.00 76.5 1.97
-Robust-Constrained (SPOC-RCD) 95.35 80.23 1.86
-Table 1: Comparison of decoding strategies across different backbone models on STL-constrained tasks.
-Our constrained decoding approaches show significant improvements in STL satisfaction and robustness while preserving task success.
+    RQ3: Does RCD achieve better task success than HCD while maintaining high STL satisfaction?
+$\bm{\phi}_{avoid}$: STL / SR (% $\uparrow$) $\bm{\phi}_{geofence}$: STL / SR (% $\uparrow$)
+Decoding SPOC Flare PoliFormer SPOC Flare PoliFormer
+Unconstrained 72.0 / 82.5 75.5 / 82.0 77.0 / 82.5 78.0 / 81.5 68.0 / 81.0 73.0 / 81.5
+Filtering 100.0 / 72.0 100.0 / 78.5 100.0 / 75.5 100.0 / 72.0 100.0 / 66.5 100.0 / 67.5
+HCD 100.0 / 72.5 100.0 / 81.0 100.0 / 78.5 100.0 / 76.5 100.0 / 67.5 100.0 / 72.5
+RCD 93.0 / 76.0 83.0 / 82.5 87.5 / 83.5 95.5 / 80.0 80.0 / 71.5 85.5 / 77.5
+Table 1: Comparison by decoding technique across models (SPOC, FLARE, PoliFormer) for specifications $\phi_{avoid}$ and $\phi_{geofence}$.
+Each cell reports STL satisfaction / success rate (%).
+Higher is better (↑).
+(a) Unconstrained v/s HCD
+(b) Unconstrained v/s RCD
+Figure 2: Qualitative comparison of decoded trajectories for a sample scene.
+Each plot shows a top down view of an overlay of trajectories starting from the white dot under the instruction “find an alarm clock”.
+The unconstrained model passes through two forbidden regions (red squares) on the way to the target object located on the table.
+In contrast, HCD (left) and RCD (right) modify the trajectories to respect STL safety specifications while still reaching the goal.
+More visualizations are available in the Appendix A.2.
 4.3 Results
-Our results are highlighted in Table 1 and Figure 2. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
+Our results are highlighted in Table 1. We also visualize sample trajectories in Figure 2 for one scene and task. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
 RQ1 – STL satisfaction.
@@
 RQ2 – Task completion.
-[…]
+Simplex-style filtering attains high satisfaction but sacrifices task success because the agent takes predefined safe actions.
+HCD shows similar behavior: although safety is maximized, success rates are consistently 5–10% lower than the baseline across models and specifications.
+However, as HCD factors in base model logits, it is able to achieve higher task satisfaction compared to Simplex-style filtering.
+In contrast, RCD preserves success rates much closer to the unconstrained level.
+For $\phi_{avoid}$, RCD achieves 82–85% success compared to 82–83% for the unconstrained controllers for Flare and PoliFormer.
+For $\phi_{geofence}$, it maintains 77–80% compared to 81–82% unconstrained for SPOC and PoliFormer.
+However, we note that RCD does not fully recover success in every case: on SPOC with $\phi_{avoid}$ and Flare with $\phi_{geofence}$, task success remains several points below the unconstrained baseline.
+Nevertheless, RCD enforces safety while avoiding the large performance penalty observed with filtering.
 RQ3 – RCD vs. HCD.
@@
 HCD enforces strict STL satisfaction that results in frequent conservatism and lower successful task completion rates.
-In contrast, RCD’s soft penalization leads to higher task success—up to 4–5 % more than HCD—while still maintaining over 93 % STL satisfaction.
+In contrast, RCD’s soft penalization leads to higher task success while still maintaining reasonable STL satisfaction.
 These results show that RCD achieves a better trade-off between safety and goal-directed behavior, especially in settings where occasional low-risk actions can lead to higher long-term rewards.
-We also perform an ablation for varying values of $\beta$ to investigate this tradeoff and provide these results in Appendix C.
-Our proposed techniques effectively enforces STL specifications during policy execution.
-HCD ensures full compliance but occasionally sacrifices task success due to strict truncation.
+Overall, Our proposed techniques effectively enforce safety STL specifications during policy execution.
+HCD ensures full compliance, but occasionally sacrifices task success due to strict truncation.
 RCD strikes a balance, offering high satisfaction rates and robust performance.
 This highlights the feasibility of combining learning-based models with formal safety constraints.
+4.4 Ablation studies
+Does inaccurate dynamics modeling substantially reduce STL satisfaction?
+In this work, we assume a simple unicycle dynamics model due it’s generalization capability for diverse robotic platforms.
+Although this represents a high-level abstraction of true dynamics, such modeling simplifications are standard in the formally assured robot safety literature [28].
+However, both RCD and HCD depend on this assumption and inaccurate modeling can impact STL satisfaction.
+To evaluate the impact of inaccurate dynamics modeling, we conducted an ablation in which we inject gaussian perturbations into the dynamics (0.01 m per step translational noise i.e.
+5% of nominal forward step, 1 ° per step rotational noise i.e.
+3.3% of yaw step) for both HCD and RCD.
+Our results are visualized in Figure 3.
+Across all three base models, the drop in STL satisfaction rates from baseline to noisy dynamics is relatively small.
+We observe that SafeDec shows graceful degradation under significant per-step disturbances.
+(a) Ablation with noisy dynamics
+(b) Ablation on varying values of $\beta$ for RCD
+Figure 3: Ablation studies. (a) STL satisfaction (%) for HCD and RCD under baseline vs noisy dynamics across base models. (b) Effect of $\beta$ on success rate and safety satisfaction.
+How do varying values of $\beta$ affect success rate and STL satisfaction in RCD?
+As highlighted in section 3.3, RCD uses a hyperparameter $\beta$ that modulates the trade-off between specification adherence and the original task objective.
+The value of $\beta$ affects the relative weighting between robustness and base logits.
+To investigate the impact of $\beta$ on the success rate and STL satisfaction, we performed an ablation with varying values for $\beta$ for Flare and PoliFormer.
+We observe that as $\beta$ increases for PoliFormer, both STL satisfaction and success rate improve in tandem until $\beta$ = 10, suggesting that moderate regularization can actually aid policy execution.
+Beyond this, STL satisfaction continues to improve but at the cost of lower success rates.
+For Flare, larger $\beta$ values improve STL satisfaction but reduce success rates.
+These results highlight that the influence of $\beta$ is model-dependent but in general demonstrate that SafeDec provides a tunable mechanism to balance safety and performance objectives.
 5 Related Work
-Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [24], safe reinforcement learning [25], and temporal logic-based shielding approaches [26].
+Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [30], safe reinforcement learning [31], temporal logic-based shielding and synthesis approaches [32, 33, 34].
 Recently, with the advent of vision language action models and their impressive generalizable capabilities for manipulation, navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models.
-Although classical methods offer formal guarantees, they require pretraining/fine-tuning stage interventions, which can be restrictive.
+Although classical methods offer formal guarantees, they either require pretraining/fine-tuning stage interventions or designing a new classical controller for each safety specification, which can be restrictive.
 For example, SafeVLA [35] fine-tunes pre-trained foundation models with task-specific safety costs, achieving strong performance in Safety-CHORES tasks.
@@
 In contrast, ASIMOV [36] explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees.
-Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model and incurring negligible computational overhead.
+Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model.
 The closest to our work is SELP [37] that proposes LTL-constrained decoding for language model-based plan generation.
-However, since LTL does not possess quantiative semantics, they are restricted to hard masking-based approach.
+However, SELP is unsuitable for STL because its Boolean predicate-based LTL cannot encode numeric bounds (e.g.
+$x–x_{goal}||<0.1m$) and does not possess quantitative semantics, which is crucial for ranking actions.
+These techniques are also tailored to high-level plans from LLMs, not to the per-step low-level actions generated online by policies.
 6 Limitations and Future Work
-In this work, we introduce a constrained decoding framework for enforcing STL specifications for robotic foundation models.
+In this work, we introduce a constrained decoding framework for enforcing safety specifications for large transformer based robot policies.
 Our approach enables runtime adaptation to novel safety specifications without retraining.
@@
 Although this is a common situation for safety critical deployment contexts like aerial robotics [38, 39], these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation.
-We hope to remedy this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic (ETL) [32] and using Large Language Models for generating high level specifications automatically [33].
-Second, our approach also assumes access to a predictive model to evaluate the impact of actions on future trajectories.
-This can limit applicability to settings where accurate dynamics models are unavailable.
-However, it is possible to mitigate this via learned dynamics models [34] or world models proposed in [35, 36].
-In addition, while our experiments are conducted in simulation, the underlying foundation model (SPOC) has been successfully deployed on real robots with minimal sim-to-real degradation [2].
-Since our constrained decoding operates primarily at inference time and leaves the base model unchanged, the observed improvements in STL satisfaction and task success are expected to transfer over to the real world.
-We plan to rigorously test this by deploying SpecDec on the same physical robot platform as [2].
-Appendix
-Appendix A Quantiative Semantics of STL
+We plan to overcome this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic [40] and using large language models for generating high level specifications automatically [41].
+Second, our approach also assumes access to an approximate dynamics model to evaluate the impact of actions on future trajectories.
+While a common assumption for provably safe robotics [28], This can limit applicability of our approach.
+However, it is possible to mitigate this via learned dynamics models such as MoSim [42] or world models proposed in Zhou et al.
+[43], Micheli et al.
+[44] and we plan to explore this in future work.
+Appendix A Appendix
+A.1 Quantitative Semantics of STL
 Given a signal $s_{t}$ representing a signal starting at time t, the quantitative semantics of satisfaction of $s_{t}\models\phi$ are defined inductively as follows:
@@
 $\rho(s_{t},\text{G}_{[a,b]}(\varphi))$ $=\underset{t^{\prime}\in[t+a,t+b]}{\min}\rho(s_{t}^{\prime},\varphi)$
-Appendix B HCD proof
-We rewrite the theorem here for easier comprehension.
-Let $\pi_{\text{HCD}}$ denote the policy induced by hard-constrained decoding with respect to an STL specification $\varphi$, and let $f$ be a deterministic dynamics model used during decoding.
-Then the generated trajectory satisfies the specification:
-$\{(x_{t+1},\hat{a}_{t+1}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
-where each action $\hat{a}_{t+k}\sim\pi_{\text{HCD}}(\cdot\mid x_{t+k-1})$, and $x_{t+k}=f(x_{t+k-1},\hat{a}_{t+k})\quad\text{for all }k.$
-Proof Sketch.
-[…]
-We proceed by induction over time.
-At step $t+1$, the masking ensures that the selected action $a_{t+1}$ is such that $(x_{t+1},a_{t+1})\models\varphi$, where $x_{t+1}=f(x_{t},a_{t+1})$.
-Assuming that the sequence $\{(x_{t+1},a_{t+1}),\dots,(x_{t+k-1},a_{t+k-1})\}$ satisfies $\varphi$, we show that $(x_{t+k},a_{t+k})$ will also satisfy $\varphi$.
-Since actions are masked based on predicted rollouts using $f$, and $f$ is exact, the only permitted actions at step $t+k$ are those that extend the sequence without violating $\varphi$.
-By induction, the entire sequence satisfies $\varphi$.
-∎
-Appendix C Robustness weighting ablation
-Figure 3: Success rate and STL satisfaction rate across different values of $\beta$ for RCD, which controls the importance of STL spcification during decoding.
-As $\beta$ increases, STL satisfaction improves while task success shows a slight tradeoff beyond moderate values.
-To evaluate the impact of relative weighting between robustness and base logits, we plotted the success rate and STL satisfaction rate across varying values for $\beta$.
-As shown in Figure 3, increasing the robustness threshold initially improves both task success and specification satisfaction, with success rates peaking around a moderate threshold (5–10).
-However, at higher thresholds (20 and 25), success rates begin to decline slightly, while STL satisfaction remains high.
-This trend highlights a trade-off introduced by weighting robustness.
-Higher robustness weighting leads to better STL satisfaction at the cost of hindering overall task completion as logits for feasible actions are shifted heavily by the robustness values.
-These results demonstrate the importance of carefully balancing robustness weighting to maintain both task success and STL satisfaction.
+A.2 Visualizations
+To complement our quantitative results, we provide numerous trajectory plots from evaluations across a diverse set of procedurally generated indoor environments.
+Figure 4 illustrates representative top-down visualizations of trajectories induced by SafeDec .Each plot shows the agent’s starting point and the resulting path under constrained decoding, with red circles marking target objects, green boxes denoting forbidden regions, and orange paths depicting the safe trajectories generated by our method.
+Together with our quantitative analysis, these qualitative results illustrate the performance of SafeDec across environments and tasks.
+Figure 4: Top-down views across indoor environments with SafeDec induced trajectories Red circles mark target objects, green boxes are avoid regions, and orange paths show safe trajectories under constrained decoding.

### v4

@@
-Constrained Decoding for Robotics Foundation Models
+SafeDec: Constrained Decoding for Safe Autoregressive Generalist Robot Navigation Policies
 Abstract
-Recent advances in the development of robotic foundation models have led to promising end-to-end and general-purpose capabilities in robotic systems.
-Trained on vast datasets of simulated and real-world trajectories, these models map multimodal observations directly to action sequences for physical execution.
+Recent advances in end-to-end, multi-task robot policies based on transformer models have demonstrated impressive generalization to real-world embodied navigation tasks.
+Trained on vast datasets of simulated and real-world trajectories, these policies map multimodal observations directly to action sequences for physical execution.
 Despite promising real-world capabilities, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness.
-We address this gap by introducing SafeDec, a constrained decoding framework for autoregressive, robot foundation models that enforces invariant safety specifications on candidate action trajectories.
-Task-specific safety rules are expressed as Signal Temporal Logic (STL) formulas and are enforced at inference time with minimal overhead.
-Our method ensures that generated actions provably satisfy STL specifications under assumed dynamics at runtime without retraining , while remaining agnostic of the underlying policy.
-We evaluate SafeDec on tasks from the CHORES benchmark for state-of-the-art generalist policies (e.g., SPOC, Flare, PoliFormer) across hundreds of procedurally generated environments and show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action generation.
+We address this gap by introducing SafeDec, a constrained decoding framework for autoregressive, transformer-based robot navigation policies that enforces safety specifications expressed as Signal Temporal Logic (STL) formulas.
+Our method ensures that generated actions provably satisfy STL specifications under assumed dynamics at runtime without retraining while remaining agnostic of the underlying policy.
+We evaluate SafeDec on tasks from the CHORES benchmark for state-of-the-art embodied navigation policies across hundreds of procedurally generated environments and show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action generation.
 Videos are available at constrained-robot-fms.github.io.
+Machine Learning, Safe Robotics, Constrained Decoding, Robot Navigation
 1 Introduction
-Recent advances in developing Robotics Foundation Models (RFM) have enabled general-purpose policies that map multi-modal inputs such as RGB images, natural language instructions, and proprioceptive inputs to action sequences Hu et al.
-[1].
-Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3], Flare [4] and OpenVLA [5] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts.
-However, these models are primarily data-driven and lack any explicit notion of safety.
-Although these models may implicitly exhibit safety-related behaviors depending on the patterns in their training data, there is no formal guarantee that models will consistently behave safely in all situations.
-This serves as a limiting factor for deploying these foundation models in the physical world where rule compliance and regulatory safety rule adherence are crucial.
-Formal specifications have long been used to specify safety requirements for robotic deployments [6, 7].
-Temporal logics (TL) [8]  can capture safety constraints on robot behavior, such as “remain within the permitted region zones and avoid dangerous obstacles”.
-Although TL has seen success in classical robotic planning for safety constraint satisfaction, its use for enforcing safety for large transformer-based robot policies remains limited.
-Additionally, retraining or fine-tuning these large pre-trained models to directly embed temporal logic specification is challenging [9].
-First, retraining models is a costly endeavor in terms of computational resources and data requirements.
+Recent advances in developing large transformer-based models for embodied navigation have enabled general-purpose policies that map multi-modal inputs such as RGB images, natural language instructions, and proprioceptive inputs to action sequences (Hu et al., 2023).
+For example, Shortest Path Oracle Clone (SPOC) (Ehsani et al., 2024), PoliFormer (Zeng et al., 2025), Flare (Hu et al., 2025), and LEO (Huang et al., 2024) exhibit impressive zero-shot generalization for embodied navigation tasks such as open vocabulary object-goal navigation (“find a mug”) and attribute-conditioned navigation (“locate the chair closest to the refrigerator in the kitchen”).
+These models have also shown promising transfer for robot navigation in real-world settings (Zeng et al., 2025; Huang et al., 2024).
+However, these policies are primarily data-driven and lack any explicit notion of safety.
+Although policies may implicitly exhibit safety-related behaviors depending on patterns in their training data, there is no formal guarantee that they will consistently behave safely in all situations.
+This limitation hinders the deployment of such policies in the physical world where adherence to explicit safety rules and regulatory constraints is essential.
+Formal specifications have long been used to specify contextual safety requirements for robotic deployments (Menghi et al., 2019; Farrell et al., 2018).
+Specifically, temporal logics (TL) (Pnueli, 1977)  provide a structured language for specifying rich safety constraints on robot behavior, ranging from simple invariants (e.g., remaining within permitted operational zones or avoiding dangerous obstacles) to contextual and history-dependent specifications.
+For example, a robot may be required to avoid certain regions until visiting another region (temporal ordering), permanently avoid a region after entering a restricted area (conditional safety) or never return to a region once it has exited.
+Such specifications naturally capture real-world operational rules that depend not only on the robot’s current state, but also on what situations the robot has previously encountered.
+Although TL has seen success in classical robotic planning and reinforcement learning for safety constraint satisfaction, its use for enforcing safety for large transformer-based robot policies remains limited (Manganaris et al., 2026).
+Additionally, retraining or fine-tuning these large pre-trained policies to directly embed such non-Markovian TL safety specifications is challenging (Kapoor et al., 2024).
+First, retraining policies is a costly endeavor in terms of computational resources and data requirements.
 Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of safety constraints through training alone.
-Hence, there is a pressing need for methods that can enforce safety specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
-In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time [10, 11, 12].
+Hence, there is a pressing need for methods that can reliably enforce safety specifications at inference time without policy retraining.
+In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time (Willard and Louf, 2023; Beurer-Kellner et al., 2023; AI, 2023).
 These approaches typically mask out tokens that violate a syntactic constraint defined over token sequences.
-For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns [10, 11].
-Inspired by this line of work, we extend the paradigm of constrained decoding to enforce safety constraints over action trajectories in dynamical systems and propose safety specification aligned decoding (SafeDec) for transformer based policies that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [13] specifications under assumed dynamics.
+For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns (Willard and Louf, 2023; Beurer-Kellner et al., 2023).
+Inspired by this line of work, we extend the paradigm of constrained decoding to enforce contextual safety constraints over action trajectories in dynamical systems.
+We propose safety specification aligned decoding (SafeDec) for transformer based policies that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) (Maler and Nickovic, 2004) specifications under assumed dynamics.
 Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction.
 This conditioning is critical because it steers the model toward generating safety specification satisfying actions rather than relying on post hoc rejection.
-SafeDec reduces risk of infeasible outputs while preserving the original action distribution of the model.
-To enforce such specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations.
+To enforce safety specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations.
 Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to an approximate dynamics model to predict future states.
-To efficiently evaluate STL specifications at inference time, we use a high-performance computational graph based library STLCG++ [14].
-To the best of our knowledge, this is the first work to enforce formal safety on transformer-based robotic policies at inference time using constrained decoding.
-Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for transformer-based policies under safety STL specifications (Section 3.1).
-Second, we propose an inference-time technique that reweights or masks candidate actions using STL satisfaction scores in (Section 3.2 and 3.3).
-Finally, we demonstrate the effectiveness of our method on state-of-the-art object navigation models without modifying model parameters (Section 4).
+In this work, we focus on safety enforcement for embodied navigation policies in indoor environments.
+We also focus on STL safety (Alpern and Schneider, 1987) specifications without liveness operators, since such specifications are sufficient to model a wide range of safety constraints encountered in robot learning (He et al., 2024; Yun et al., 2025; Kapoor et al., 2025b; Zhao et al., 2024).
+To the best of our knowledge, this is the first work to enforce STL safety specifications on transformer-based navigation policies at inference time using constrained decoding.
 2 Preliminaries
 2.1 Signal Temporal Logic
-Signal Temporal Logic (STL) is an expressive framework for defining properties and reasoning over continuous time real valued signals [15].
+Signal Temporal Logic (STL) is an expressive framework for defining properties and reasoning over continuous time real-valued (Maler and Nickovic, 2004) .
 Formally, $(\mathbf{s},t)\models\phi$ denotes that a signal $\mathbf{s}$ satisfies the STL formula $\phi$ at time $t$.
@@
 Note that with slight abuse of notation, $\mu$ represents both the predicate and a function of the trajectory $\mathbf{s}(t)$.
-Any STL formula consists of Boolean and temporal operations on these predicates, and the syntax of STL formulas is defined recursively as follows:
+Any STL formula consists of Boolean and temporal operations on predicates, and the syntax of STL formulas is defined recursively as follows:
 $\phi:=\mu~|~\neg\mu~|~\phi\land\psi~|~\phi\lor\psi~|~\mathbf{G}_{[a,b]}~\psi~|~\mathbf{F}_{[a,b]}~\psi~|~\phi~\mathbf{U}_{[a,b]}~\psi$
@@
 $s_{t}\models\text{G}_{[a,b]}(\varphi)$ $\iff\forall t^{\prime}\in[t+a,t+b]\text{ s.t. }s_{t^{\prime}}\models\varphi$
-Apart from the Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric indicating robustness, i.e., the strength of satisfaction or violation.
-For the sake of brevity, the definition of robustness is provided in Appendix A.1.
+Apart from Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric called robustness, i.e., the strength of satisfaction or violation.
+The definition of robustness is provided in Appendix A.1.
 2.2 Constrained Decoding in Transformers
 A large variety of autoregressive transformer-based models generate final outputs by producing a probability distribution over the model vocabulary at each timestep.
-This distribution is generated by performing a softmax operation over the model’s last hidden layer.
 Then, through the process of decoding, tokens are selected to maximize the overall likelihood of an output sequence.
 In standard decoding, this maximization can be performed by either greedily selecting the most probable token at each step or by using a beam search to maintain multiple high-likelihood candidates.
-However, this often leads to degenerate output sequences that are repetitive [16].
-A common approach is to use sampling strategies like top $k$ [17], and nucleus sampling [16] that introduce stochasticity to encourage more diverse outputs.
-Constrained decoding [18] modifies this probabilistic selection by pruning invalid tokens to ensure that the generated sequences satisfy predefined constraints.
-These constraints are often syntactic, such as regular expressions, JSON formatting, or programming language grammars [19].
-There is also recent work on enforcing semantic constraints that ensure coherence of the output or alignment with specific knowledge bases [20].
+However, this often leads to degenerate output sequences that are repetitive (Holtzman et al., 2019).
+A common approach is to use sampling strategies like top $k$ (Fan et al., 2018), and nucleus sampling (Holtzman et al., 2019) that introduce stochasticity to encourage more diverse outputs.
+Constrained decoding (Hokamp and Liu, 2017) modifies this probabilistic selection by pruning invalid tokens to ensure that the generated sequences satisfy predefined constraints.
+These constraints are often syntactic, such as regular expressions, JSON formatting, or programming language grammars (Welleck et al., 2024).
+There is also recent work on enforcing semantic constraints that ensure coherence of the output or alignment with specific knowledge bases (Peyrard et al., 2024).
 Formally, constrained decoding can be seen as maximizing the probability of the output sequence subject to a constraint $\mathcal{C}$: $\arg\max_{y\in\mathcal{Y}_{\mathcal{C}}}P(y\mid x)$ where $\mathcal{Y}_{\mathcal{C}}$ is the set of sequences satisfying $\mathcal{C}$.
@@
 Figure 1: Overview of our specification aligned decoding framework.
-Given multimodal inputs (e.g., RGB images and natural language instructions), a pretrained robot foundation model (e.g., SPOC) generates candidate actions.
-These actions are filtered or reweighted by the constrained decoding technique based on STL safety constraints.
-In the figure, green markers denote target object locations, while red zones represent regions to avoid.
-In this section, we introduce a novel problem formulation for SafeDec in RFMs.
-First, we highlight the challenge in specification checking for these models in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it.
+Given multimodal observations (RGB images and language goals), a pretrained transformer-based navigation policy generates candidate actions.
+Our constrained decoder then filters or reweights these actions using robustness scores from a user-defined STL safety specification.
+Left: Original action probabilities (red) from the pretrained policy are modified by SafeDec using STL robustness: actions predicted to violate the safety specification receive reduced weights, while safer actions are boosted.
+The chosen safe action and adjusted probabilities (green) are highlighted.
+Right: An example navigation episode where the robot starts in the kitchen (white dot) and moves toward the bedroom to locate the target object (green marker middle-right), while avoiding user-defined hazardous regions (red zones) throughout the trajectory.
+In this section, we introduce a novel problem formulation for SafeDec in autoregressive transformer-based policies.
+First, we highlight the challenge in specification checking for these policies in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it.
 Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
@@
 In these setups, constraint checking can be performed in the model’s token space.
-In contrast, RFMs operate in a physical environment and constraints (captured via temporal logic) are defined over state variables.
+In contrast, transformer-based policies operate in a physical environment and constraints (captured via temporal logic) are defined over state variables.
 Since a large class of end-to-end policies solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward.
@@
 This system is controlled by a policy that selects an action $a_{t}$ at each time step based on observations and task context such as user-provided natural-language instructions or goal waypoints.
-In this work, we focus on RFMs that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions.
+In this work, we focus on robot policies that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions.
 Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$.
@@
 $\{(x_{0},\hat{a}_{0}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
-Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed actions through filtering or rejecting action sequences that violate the specification $\varphi$.
-Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low-likelihood output.
-This undermines the inductive biases learned during pretraining and leads to degenerate, brittle behaviors.
+Most of the existing techniques for specification enforcement perform post-hoc manipulation of proposed actions (after sampling from the policy) by filtering or rejecting action sequences that violate the specification $\varphi$.
+Although manipulation after sampling can ensure specification satisfaction, it can favor safe but low-likelihood actions and unintentionally suppress feasible safe behaviors that the model would have selected if safety constraints had been considered pre-sampling.
+This undermines the inductive biases learned during pretraining and lead to degenerate and brittle behaviors.
 A similar problem was highlighted when ensuring compliance with logical constraints for large language models (LLM) in Park et al.
-[21].
+(2024).
 Additionally, these models decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$.
@@
 In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques.
-Specifically, we propose SafeDec : A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
+Specifically, we propose SafeDec : A constrained decoding strategy that integrates STL specifications into the policy’s action selection process, ensuring satisfaction while remaining as close as possible to the base model distribution.
 3.2 Hard Constrained Decoding
@@
 These logits are further converted into a probability distribution using a softmax operation.
-In prior work, for structured output generation in LLMs, some invalid tokens are masked based on syntactical constraints or other criteria [19, 21].
+In prior work, for structured output generation in LLMs, some invalid tokens are masked based on syntactical constraints or other criteria (Welleck et al., 2024; Park et al., 2024).
 This is done by setting their logit value as -$\infty$ before the softmax operation is applied.
-For HCD, we use a similar approach as constrained decoding literature [19] and mask out predicted action tokens that violate our given STL specification $\varphi$ during sequential generation.
+For HCD, we use a similar approach as constrained decoding literature (Welleck et al., 2024) and mask out predicted action tokens that violate our given STL specification $\varphi$ during sequential generation.
 Formally, to enforce the STL specification $\varphi$ during sequential generation, we adjust the logits at each timestep $t+k$ as follows:
 Let $\textbf{z}_{t+k}$ denote the logits at timestep $t+k$. For each action choice $i$ at timestep $t+k$, we define:
-$z_{t+k}^{(i)}=\begin{cases}-\infty,&\text{if }\hat{x}_{t+k}^{(i)}=f(x_{t+k-1},\hat{a}_{t+k}^{(i)})\text{ violates }\varphi\\ z_{t+k}^{(i)},&\text{otherwise}\end{cases}$
-Here $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action.
+$z_{t+k}^{(i)}=\begin{cases}-\infty,&\text{if }\hat{\tau}^{(i)}_{0:t+k}=(x_{0},x_{1},..,\hat{x}_{t+k}^{(i)})\text{ violates }\varphi\\ z_{t+k}^{(i)},&\text{otherwise}\end{cases}$
+where $\hat{x}_{t+k}^{(i)}=f(\hat{x}_{t+k-1},\hat{a}_{t+k}^{(i)})$.
+Here, $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action.
 This next state is elicited using a simple dynamics model ($f$) as highlighted in the previous section.
 Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function.
+Table 1: STL specifications used in our evaluation. Each region $R_{i}$ denotes a spatial predicate over $(x,z)$.
+Spec name STL formula Description
+$\phi_{\text{avoid}}$ $\mathbf{G}\!\left(\bigwedge_{i=1}^{N}\neg R_{i}\right)$ Avoidance: the robot must always stay outside all unsafe regions.
+$\phi_{\text{ordered}}$ $\mathbf{G}(\neg R_{2})\ \vee\ (\neg R_{2}\ \mathbf{U}\ R_{1})$ Weak-until ordering: the robot must avoid region $R_{2}$ until reaching region $R_{1}$, or avoid $R_{2}$ forever.
+$\phi_{\text{no\_return}}$ $\mathbf{G}\big(R_{0}\Rightarrow\mathbf{G}(\neg R_{0}\Rightarrow\mathbf{G}\neg R_{0})\big)$ Conditional no-return: if the robot is in region $R_{0}$, then once it leaves $R_{0}$, it may never re-enter.
+$\phi_{\text{geofence}}$ $\mathbf{G}\!\left(\bigvee_{i=1}^{N}R_{i}\right)$ Geofencing: the robot must always remain within at least one allowed region.
 3.3 Robustness Constrained Decoding
@@
 A similar tradeoff was observed by Liu et al.
-[22] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity.
+(2021) when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity.
 Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness).
@@
 This is similar to the approach proposed in Liu et al.
-[22] where the next-token distributions were re-weighted based on the utility scores provided by another language model.
-Our utility scores are quantified by the robustness function( $\rho(x_{t},\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification.
+(2021) where the next-token distributions were re-weighted based on the utility scores provided by another language model.
+Our utility scores are quantified by the robustness function( $\rho(\langle x_{0},x_{1},\dots,x_{t}\rangle,\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification.
 Positive robustness values denote specification satisfaction, while negative values capture the degree of violation.
 A formal definition of robustness in line with the STL quantitative semantics is provided in Appendix A.1.
-First, we compute a robustness score for each candidate action: $r_{t+k}^{(i)}=\rho(\hat{x}_{t+k}^{(i)},\varphi)$ where $\rho(\cdot,\varphi)$ is the STL robustness metric, and $\hat{x}_{t+k}^{(i)}$ is the predicted next state under action $\hat{a}_{t+k}^{(i)}$.
+First, we compute a robustness score for each candidate action: $r_{t+k}^{(i)}=\rho(\langle x_{0},x_{1},\dots,x_{t+k-1},\hat{x}_{t+k}^{(i)}\rangle,\varphi)$ where $\rho(\cdot,\varphi)$ is the STL robustness metric, and $\hat{x}_{t+k}^{(i)}$ is the predicted next state under action $\hat{a}_{t+k}^{(i)}$.
 This robustness score $r_{t+k}^{(i)}$ quantifies how well each candidate action satisfies the specification $\varphi$.
-These scores are then converted into weights using exponential scaling: $w_{t+k,i}=\exp(\alpha\cdot r_{t+k,i})$ where $\alpha$ is a temperature parameter that adjusts the sharpness of the bias.
+These scores are then converted into weights using exponential scaling: $w_{t+k,i}=\exp(\alpha\cdot r_{t+k}^{(i)})$ where $\alpha$ is a temperature parameter that adjusts the sharpness of the bias.
 We use these weights to shift the original logits: $\tilde{z}_{t+k,i}=z_{t+k,i}+\beta\cdot w_{t+k,i}$ where $\beta$ is a hyperparameter that modulates the trade-off between specification adherence and the original task objective.
@@
 4.1 Implementational Details
-We comprehensively evaluate our constrained decoding framework on procedurally generated AI2-THOR [23] indoor scenes with diverse objects and layouts using three state-of-the-art (SOTA) RFMs: Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3] and Flare [4].
-All three are large transformer-based embodied agents trained on extensive language-conditioned robot trajectory datasets.
+We comprehensively evaluate our constrained decoding framework on procedurally generated AI2-THOR (Kolve et al., 2022) indoor scenes with diverse objects and layouts.
+AI2-THOR is a photorealistic 3D simulation environment built on the Unity engine and is a widely adopted benchmark and training environment for embodied navigation research.
+We use three state-of-the-art (SOTA) generalist robot navigation policies: Shortest Path Oracle Clone (SPOC) (Ehsani et al., 2024), PoliFormer (Zeng et al., 2025) and Flare (Hu et al., 2025).
+All three are large transformer-based policies trained on extensive language-conditioned robot trajectory datasets.
 These models achieve strong zero-shot generalization for a vast variety of navigation tasks that span open vocabulary object-goal navigation (“find a mug”), room-to-room traversal (“visit all rooms”), waypoint-based navigation (“move three meters forward and stop near the red rug”), and attribute-conditioned variants (“locate the chair closest to the refrigerator in the kitchen”).
 These models also demonstrate reliable zero-shot transfer to real-world environments, achieving robust task satisfaction.
-In addition, these models capture three different training paradigms for RFMs.
+In addition, these models capture three different training paradigms for generalist robot navigation policies.
 SPOC is trained purely with imitation learning from shortest-path rollouts.
@@
 In this work, we address safety specifications for robotics and, therefore, select those most relevant to real-world deployment.
-In particular, we focus on an important class of safety specifications called invariants, which are specifications that must be enforced at every reachable state of the system (e.g., “always avoid an unsafe region”).
-We enforce geofencing and obstacle avoidance by encoding them as invariant specifications in STL.
-Specifically, we generate random regions in the configuration space that the robot must either avoid (obstacle zones) or remain within (safe zones), and apply these constraints in real time during execution.
-The specifications used are: $\varphi_{\text{geofence}}=\mathbf{G}\!\left(\bigvee_{i=1}^{N}\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$, $\varphi_{\text{avoid}}\;=\;\mathbf{G}\!\left(\bigwedge_{i=1}^{N}\neg\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$.
-The size of the regions for $\varphi_{avoid}$ is 1 $m^{2}$.
-For $\varphi_{geofence}$, we randomly pick a subset of rooms in each house and use each chosen room’s full bounds.
-These safety specifications are common for robot learning applications [24, 25, 26, 27].
-We encode our test STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [14].
+Our specifications are highlighted in Table 1.
+These specifications are inspired by common safe navigation specifications as highlighted in Menghi et al.
+(2019) and have varying levels of complexity.
+These specifications capture state-based invariants ($\phi_{avoid},\phi_{geofence}$), history-dependent conditionals ($\phi_{no\_return}$) and temporal ordering constraints ($\phi_{ordered}$).
+Table 2: Comparison by decoding technique across models (SPOC, Flare, PoliFormer) for specifications $\phi_{avoid}$, $\phi_{ordered}$, and $\phi_{no\_return}$.
+Each cell reports STL satisfaction / success rate (%).
+Higher is better (↑).
+$\bm{\phi}_{avoid}$: STL St / SR (% $\uparrow$) $\bm{\phi}_{ordered}$: STL St / SR (% $\uparrow$) $\bm{\phi}_{no\_return}$: STL St / SR (% $\uparrow$)
+Decoding SPOC Flare PoliFormer SPOC Flare PoliFormer SPOC Flare PoliFormer
+Unconstrained 72.0 / 82.5 75.5 / 82.0 77.0 / 82.5 76.0 / 80.0 72.0 / 80.5 71.0 / 81.5 84.0 / 79.5 82.0 / 81.0 82.0 / 79.0
+Filtering 100.0 / 72.0 100.0 / 78.5 100.0 / 75.5 100.0 / 69.0 100.0 / 70.0 100.0 / 67.5 100.0 / 68.5 100.0 / 74.0 100.0 / 66.5
+HCD 100.0 / 72.5 100.0 / 81.0 100.0 / 78.5 100.0 / 70.5 100.0 / 73.0 100.0 / 72.5 100.0 / 71.0 100.0 / 76.5 100.0 / 69.5
+RCD 93.0 / 76.0 83.0 / 82.5 87.5 / 83.5 91.0 / 75.0 89.0 / 79.5 95.0 / 78.0 96.0 / 77.5 96.0 / 80.0 94.0 / 74.0
+We encode our test STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel (Kapoor et al., 2025c).
 This ensures minimal inference overhead at runtime ($10^{-5}$ s per timestep) , which is crucial for policy deployment.
-For our dynamics model, we assume a unicycle model, an approximate first-order dynamics abstraction widely used in the robotics literature for analysis and control [28].
+For our specifications, we generate random regions in the configuration space at runtime and use their coordinates as spatial predicates.
+For our dynamics model, we assume a unicycle model, an approximate first-order dynamics abstraction widely used in the robotics literature for analysis and control (Cohen et al., 2024).
 This representation captures the essential kinematics of motion in the plane and is widely used because it is applicable for diverse robotic platforms.
@@
 We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism.
-The filtering mechanism picks a default action (turning left or right in place) upon predicted violation of the safety specification, similar to the Simplex architecture [29].
+The filtering mechanism picks default actions that cannot immediately violate the specification (eg.
+turning in place) upon predicted violation of the safety specification, similar to the Simplex architecture (Sha, 2001).
 Simplex architecture is a classic scheme in which a high-performance advanced controller is continuously monitored by a provably safe but less capable backup controller.
 Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons.
+Additionally, our base policies represent actions as a discrete set of low-level action primitives (e.g., move forward/backward, lateral strafing, rotation, end, and a few other discrete navigation actions exposed by each policy).
+At each timestep, the policy outputs a categorical distribution over this shared action set.
 We evaluate performance using two main metrics: STL Satisfaction Rate (STL St), defined as the proportion of trajectories that satisfy the specified STL formula, and Task Success Rate (SR), which measures standard task success.
-The three main research questions we investigated in this paper:
-  1. 1.
-    RQ1: Do HCD and RCD provide higher STL satisfaction than the unconstrained baselines?
-  2. 2.
-    RQ2: Do HCD and RCD preserve task success rates comparable to the unconstrained baselines?
-  3. 3.
-    RQ3: Does RCD achieve better task success than HCD while maintaining high STL satisfaction?
-$\bm{\phi}_{avoid}$: STL / SR (% $\uparrow$) $\bm{\phi}_{geofence}$: STL / SR (% $\uparrow$)
-Decoding SPOC Flare PoliFormer SPOC Flare PoliFormer
-Unconstrained 72.0 / 82.5 75.5 / 82.0 77.0 / 82.5 78.0 / 81.5 68.0 / 81.0 73.0 / 81.5
-Filtering 100.0 / 72.0 100.0 / 78.5 100.0 / 75.5 100.0 / 72.0 100.0 / 66.5 100.0 / 67.5
-HCD 100.0 / 72.5 100.0 / 81.0 100.0 / 78.5 100.0 / 76.5 100.0 / 67.5 100.0 / 72.5
-RCD 93.0 / 76.0 83.0 / 82.5 87.5 / 83.5 95.5 / 80.0 80.0 / 71.5 85.5 / 77.5
-Table 1: Comparison by decoding technique across models (SPOC, FLARE, PoliFormer) for specifications $\phi_{avoid}$ and $\phi_{geofence}$.
-Each cell reports STL satisfaction / success rate (%).
-Higher is better (↑).
-(a) Unconstrained v/s HCD
-(b) Unconstrained v/s RCD
-Figure 2: Qualitative comparison of decoded trajectories for a sample scene.
-Each plot shows a top down view of an overlay of trajectories starting from the white dot under the instruction “find an alarm clock”.
-The unconstrained model passes through two forbidden regions (red squares) on the way to the target object located on the table.
-In contrast, HCD (left) and RCD (right) modify the trajectories to respect STL safety specifications while still reaching the goal.
-More visualizations are available in the Appendix A.2.
 4.3 Results
-Our results are highlighted in Table 1. We also visualize sample trajectories in Figure 2 for one scene and task. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
-RQ1 – STL satisfaction.
-[…]
-RQ2 – Task completion.
-Simplex-style filtering attains high satisfaction but sacrifices task success because the agent takes predefined safe actions.
-HCD shows similar behavior: although safety is maximized, success rates are consistently 5–10% lower than the baseline across models and specifications.
-However, as HCD factors in base model logits, it is able to achieve higher task satisfaction compared to Simplex-style filtering.
-In contrast, RCD preserves success rates much closer to the unconstrained level.
-For $\phi_{avoid}$, RCD achieves 82–85% success compared to 82–83% for the unconstrained controllers for Flare and PoliFormer.
-For $\phi_{geofence}$, it maintains 77–80% compared to 81–82% unconstrained for SPOC and PoliFormer.
-However, we note that RCD does not fully recover success in every case: on SPOC with $\phi_{avoid}$ and Flare with $\phi_{geofence}$, task success remains several points below the unconstrained baseline.
+Our results are highlighted in Table 2.
+For brevity, we provide the results for $\phi_{geofence}$ in 4 since it is syntactically similar to $\phi_{avoid}$.
+We also visualize sample trajectories in Figure 2 for one scene and task.
+Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
+Do HCD and RCD provide higher STL satisfaction than the unconstrained baselines?
+Both HCD and RCD consistently improve STL satisfaction relative to the unconstrained baselines across all models and specifications.
+All policies with RCD show significant gains, with improvements ranging from 8-24 percentage points.
+HCD achieves perfect STL satisfaction, with improvements ranging from 16-29 percentage points.
+We observe that the Simplex-style filtering baseline achieves similar STL satisfaction rate as HCD.
+This parity is expected, as both methods block any action predicted to violate the specification.
+Figure 2: Qualitative comparison (Unconstrained vs.
+HCD).
+Top-down visualization of trajectories for a sample scene under the instruction “find the sofa”.
+The unconstrained policy passes through forbidden regions (red squares), while Hard-Constrained Decoding (HCD) enforces STL safety specifications and generates trajectories that avoid unsafe regions while still reaching the sofa (center-left).
+Additional visualizations are provided in Appendix A.2.
+Which safety enforcement technique achieves task success rates comparable to the unconstrained baselines?
+As highlighted in Table 2, Simplex-style filtering attains high STL satisfaction but sacrifices task success because the agent takes predefined safe actions.
+HCD shows similar behavior: although safety is maximized, success rates are consistently lower than the unconstrained baseline across models and specifications.
+However, as HCD factors in base model logits, it is able to achieve higher task satisfaction compared to Simplex-style filtering with improvements ranging from 0.5-3 percentage points.
+In contrast, RCD preserves success rates much closer to the unconstrained level with success-rate drops of 0.5–6.5 percentage points (smaller drop is better).
+However, we note that RCD does not fully recover success for certain models and specifications, task success remains several points below the unconstrained baseline.
 Nevertheless, RCD enforces safety while avoiding the large performance penalty observed with filtering.
-RQ3 – RCD vs. HCD.
+Does RCD achieve better task success than HCD while maintaining high STL satisfaction?
 While both HCD and RCD improve safety over the unconstrained baseline, they differ in how they balance constraint satisfaction with task success.
@@
 RCD strikes a balance, offering high satisfaction rates and robust performance.
-This highlights the feasibility of combining learning-based models with formal safety constraints.
+4.4 Comparison with safe reinforcement learning baseline
+Table 3: Comparison with SafeVLA on the SafetyChores benchmark. Lower cost indicates fewer safety violations, while higher success denotes better task completion.
+Method Cost $\downarrow$ SR (%) $\uparrow$
+ISA (SafeVLA) 0.205 86.5
+ISA + HCD 0.015 82.0
+ISA + RCD 0.060 86.0
+FLARE 0.192 82.0
+FLARE + HCD 0.115 84.0
+FLARE + RCD 0.155 79.0
+We additionally compare SafeDec with SafeVLA (Zhang et al., 2025), a safe reinforcement learning method that proposes an integrated safety approach (ISA) using explicit safety costs.
+To enable a fair comparison, we focus on the fragile-object and dangerous-object specifications in the SafetyChores benchmark (Zhang et al., 2025), which closely align with our setting, and construct corresponding STL specifications for evaluation.
+We report the composite safety cost defined in SafeVLA, which directly reflects the degree of safety violations across both object classes, along with task success.
+Our results are highlighted in Table 3.
+Our findings indicate that both HCD and RCD substantially reduce safety violations relative to their respective vanilla baselines while largely preserving task success.
+In particular, integrating HCD with ISA (SafeVLA) achieves an order-of-magnitude reduction in safety cost (0.205 → 0.015), demonstrating effective enforcement of fragile and dangerous object specifications with minimal impact on task performance.
+We also find that RCD+ISA preserves task success close to vanilla ISA while reducing safety violations.
+Our results show that SafeDec, despite being a training-free method, preserves task satisfaction close to the safeRL baseline (ISA) for our specifications.
+Additionally, combining SafeDec with the ISA policy leads to further reduction in safety costs, highlighting its complementary benefits to data-driven safe learning.
 4.5 Ablation studies
@@
 In this work, we assume a simple unicycle dynamics model due it’s generalization capability for diverse robotic platforms.
-Although this represents a high-level abstraction of true dynamics, such modeling simplifications are standard in the formally assured robot safety literature [28].
+Although this represents a high-level abstraction of true dynamics, such modeling simplifications are standard in the formally assured robot safety literature (Cohen et al., 2024).
 However, both RCD and HCD depend on this assumption and inaccurate modeling can impact STL satisfaction.
@@
 3.3% of yaw step) for both HCD and RCD.
-Our results are visualized in Figure 3.
+Our results are visualized in Figure 3(a).
 Across all three base models, the drop in STL satisfaction rates from baseline to noisy dynamics is relatively small.
@@
 To investigate the impact of $\beta$ on the success rate and STL satisfaction, we performed an ablation with varying values for $\beta$ for Flare and PoliFormer.
-We observe that as $\beta$ increases for PoliFormer, both STL satisfaction and success rate improve in tandem until $\beta$ = 10, suggesting that moderate regularization can actually aid policy execution.
+As visualized in Figure 3(b), We observe that as $\beta$ increases for PoliFormer, both STL satisfaction and success rate improve in tandem until $\beta$ = 10, suggesting that moderate regularization can actually aid policy execution.
 Beyond this, STL satisfaction continues to improve but at the cost of lower success rates.
@@
 5 Related Work
-Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [30], safe reinforcement learning [31], temporal logic-based shielding and synthesis approaches [32, 33, 34].
-Recently, with the advent of vision language action models and their impressive generalizable capabilities for manipulation, navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models.
-Although classical methods offer formal guarantees, they either require pretraining/fine-tuning stage interventions or designing a new classical controller for each safety specification, which can be restrictive.
-For example, SafeVLA [35] fine-tunes pre-trained foundation models with task-specific safety costs, achieving strong performance in Safety-CHORES tasks.
-However, the safety specification is expected to be embedded in the training data and loss, meaning the model cannot generalize to new safety constraints at test time.
-In contrast, ASIMOV [36] explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees.
+Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) (Ames et al., 2019), safe reinforcement learning (Gu et al., 2024), and temporal logic-based shielding approaches (Alshiekh et al., 2017).
+Recently, with the advent of vision language action models and their impressive generalizable capabilities for embodied navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models.
+Although classical methods offer formal guarantees, they either require training stage interventions or designing a new classical controller for each safety specification, which can be restrictive.
+For example, SafeVLA (Zhang et al., 2025) uses safeRL with task-specific safety costs, achieving strong performance in SafetyCHORES tasks.
+However, the safety specification is expected to be embedded in the training data and loss, meaning that the model cannot generalize to new safety constraints at test time.
+In contrast, ASIMOV (Sermanet et al., 2025) explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees.
 Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model.
-The closest to our work is SELP [37] that proposes LTL-constrained decoding for language model-based plan generation.
+There has been some work on logic constrained decoding such as SELP (Wu et al., 2025) that proposes LTL-constrained decoding for language model-based plan generation.
 However, SELP is unsuitable for STL because its Boolean predicate-based LTL cannot encode numeric bounds (e.g.
 $x–x_{goal}||<0.1m$) and does not possess quantitative semantics, which is crucial for ranking actions.
-These techniques are also tailored to high-level plans from LLMs, not to the per-step low-level actions generated online by policies.
+Moreover, SELP operates at the symbolic planning level, constraining high-level plans produced by LLMs offline, whereas our method enforces safety over per-step low-level actions generated online by embodied policies.
 6 Limitations and Future Work
-In this work, we introduce a constrained decoding framework for enforcing safety specifications for large transformer based robot policies.
+In this work, we introduce a constrained decoding framework for enforcing safety specifications for large transformer based robot navigation policies.
 Our approach enables runtime adaptation to novel safety specifications without retraining.
-Through experiments across multiple simulated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates.
+Through experiments across multiple procedurally generated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates.
 Our approach makes two critical assumptions that can be a limiting factor.
 First, we assume access to specifications that are defined over the state space and that these specifications are generated by roboticists.
-Although this is a common situation for safety critical deployment contexts like aerial robotics [38, 39], these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation.
-We plan to overcome this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic [40] and using large language models for generating high level specifications automatically [41].
+Although this is a common situation for safety critical deployment contexts like aerial robotics (Aloor et al., 2023; Luckcuck et al., 2019), these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation.
+We plan to overcome this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic (Kapoor et al., 2025a) and using large language models for generating high level specifications automatically (Li et al., 2025).
 Second, our approach also assumes access to an approximate dynamics model to evaluate the impact of actions on future trajectories.
-While a common assumption for provably safe robotics [28], This can limit applicability of our approach.
-However, it is possible to mitigate this via learned dynamics models such as MoSim [42] or world models proposed in Zhou et al.
-[43], Micheli et al.
-[44] and we plan to explore this in future work.
+While a common assumption for provably safe robotics (Cohen et al., 2024), this can limit applicability of our approach.
+However, it is possible to mitigate this via learned dynamics models such as MoSim (Hao et al., 2025) or world models proposed in Zhou et al.
+(2025); Micheli et al.
+(2023) and we plan to explore this in future work.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Appendix
@@
 To complement our quantitative results, we provide numerous trajectory plots from evaluations across a diverse set of procedurally generated indoor environments.
-Figure 4 illustrates representative top-down visualizations of trajectories induced by SafeDec .Each plot shows the agent’s starting point and the resulting path under constrained decoding, with red circles marking target objects, green boxes denoting forbidden regions, and orange paths depicting the safe trajectories generated by our method.
+Figure 4 illustrates representative top-down visualizations of trajectories induced by SafeDec .
+Each plot shows the agent’s starting point and the resulting path under constrained decoding, with red circles marking target objects, green boxes denoting forbidden regions, and orange paths depicting the safe trajectories generated by our method.
 Together with our quantitative analysis, these qualitative results illustrate the performance of SafeDec across environments and tasks.
 Figure 4: Top-down views across indoor environments with SafeDec induced trajectories Red circles mark target objects, green boxes are avoid regions, and orange paths show safe trajectories under constrained decoding.
+A.3 Results for $\phi_{geofence}$
+Table 4: Performance comparison for the geofence specification $\phi_{geofence}$ across decoding techniques and base models.
+Each cell reports STL satisfaction / success rate (%).
+Higher is better (↑).
+$\bm{\phi}_{geofence}$: STL St / SR (% $\uparrow$)
+Decoding SPOC Flare PoliFormer
+Unconstrained 78.0 / 81.5 68.0 / 81.0 73.0 / 81.5
+Filtering 100.0 / 72.0 100.0 / 66.5 100.0 / 67.5
+HCD 100.0 / 76.5 100.0 / 67.5 100.0 / 72.5
+RCD 95.5 / 80.0 80.0 / 71.5 85.5 / 77.5